• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name vasulk

    Log entry time 08:07:37 on February12,2003

    Entry number 96124

    keyword=Owl Shift Update

    04:48 MCC is getting ready to send CW beam. They also need to reboot an IOC for
    the BPM's.

    04:55 MCC will send low current CW beam to the Hall.
    05:08 Move target to BeO to check spot and position.
    05:15 Raster appears a little too big, and target OTR is saturated
    MCC is working on the OTR

    05:25 OTR is working, adjusted raster a little, still a little larger, but acceptable
    05:26 Move to Polarized Helium-3 position

    05:44 Begin Kin 6.3.3, BHWP IN, 0 degrees at 14.7 microAmps
    Left and Right arm at 32 degrees, and P0 = 2.379 GeV/c

    05:50-6:00 Frequent beam trips at 14.7 microAmps
    smooth running until end of shift

    Run Summary

    Start Kin 6.3.3, BHWP IN, 0 degrees

    run # 20947/1947: from 05:44 to 06:50
    run # 20948/1948: from 06:52 to 07:52
    run # 20949/1949: from 07:52 to next shift